OpenStudy (anonymous):

The mean of the set of data below rounded to the nearest tenth is . 11, 16, 2, 2, 2, 5, 7, 7, 8

jimthompson5910 (jim_thompson5910):

Add up the numbers Then divide that result by 9 (since there are 9 numbers)

jimthompson5910 (jim_thompson5910):

Mean = (Sum of the values)/(Number of values) Mean = (11+16+2+2+2+5+7+7+8)/9 Mean = 60/9 Mean = 6.67 which rounds to 6.7